# Susceptibilities and Taylor coefficients of magnetic QCD from perturbation theory

### v1

@@ IV Summary and outlook
In this paper, we computed the coefficients $c_{2}(T,B)$ and $c_{4}(T,B)$ of the Taylor expansion for the pressure in powers of $\mu_{B}/T$ within perturbative QCD at finite temperature and density, and in the presence of very high magnetic fields, up to two loops with physical quark masses. We also calculated the excess of pressure, baryon density and baryon number susceptibility as functions of $\hat{\mu}_{B}$, as well as susceptibilities as functions of the temperature in the $\{\mu_{B},\mu_{Q},\mu_{S}\}$ basis. Since we adopt the lowest-Landau level approximation in order to obtain analytic results and more control on qualitative aspects, the region of validity for our framework is restricted to $m_{s}\ll T\ll\sqrt{eB}$, where $m_{s}$ is the strange quark mass, $e$ is the fundamental electric charge, $T$ is the temperature, and $B$ is the magnetic field strength. Even though current lattice results do not overlap with its region of validity, we found that perturbative results are compatible with those obtained on the lattice for the largest temperatures probed [56].
The magnetic perturbative series exhibits mostly narrow error bands, which originate in the truncation of the perturbative series and the mass scale that emerges as a consequence, giving more robustness to the overlap in $c_{4}$ and the good tendency in $c_{2}$ when compared to lattice data in their temperature dependence. As for their magnetic field dependence, we provide predictions to $c_{2}(T,B)$ to larger fields at high temperature, $c_{4}(T,B)$ being vanishingly small. Those are qualitatively in line with results for three smaller values of the field shown in Ref. [56]. Our findings for the excess of pressure, baryon density and baryon number susceptibility are also in good agreement with lattice data for small values of $\hat{\mu}_{B}$. Even for the smaller intensity of magnetic field considered in the lattice simulations of Ref. [76], our results for the susceptibilities $\chi_{BB}$, $\chi_{QQ}$ and $\chi_{SS}$ are reasonable when compared to their high end in temperatures.

### v2

@@ IV Summary and outlook
In this paper, we computed the coefficients $c_{2}(T,B)$ and $c_{4}(T,B)$ of the Taylor expansion for the pressure in powers of $\mu_{B}/T$ within perturbative QCD at finite temperature and density, and in the presence of very high magnetic fields, up to two loops with physical quark masses. We also calculated the excess of pressure, baryon density and baryon number susceptibility as functions of $\hat{\mu}_{B}$, as well as susceptibilities as functions of the temperature in the $\{\mu_{B},\mu_{Q},\mu_{S}\}$ basis. Since we adopt the lowest-Landau level approximation in order to obtain analytic results and more control on qualitative aspects, the region of validity for our framework is restricted to $m_{s}\ll T\ll\sqrt{eB}$, where $m_{s}$ is the strange quark mass, $e$ is the fundamental electric charge, $T$ is the temperature, and $B$ is the magnetic field strength. Even though current lattice results do not overlap with its region of validity, we found that perturbative results are compatible with those obtained on the lattice for the largest temperatures probed Astrakhantsev et al. [2024].
The magnetic perturbative series exhibits mostly narrow error bands, which originate in the truncation of the perturbative series and the mass scale that emerges as a consequence, giving more robustness to the overlap in $c_{4}$ and the good tendency in $c_{2}$ when compared to lattice data in their temperature dependence. As for their magnetic field dependence, we provide predictions to $c_{2}(T,B)$ to larger fields at high temperature, $c_{4}(T,B)$ being vanishingly small. Those are qualitatively in line with results for three smaller values of the field shown in Ref. Astrakhantsev et al. [2024]. Our findings for the excess of pressure, baryon density and baryon number susceptibility are also in good agreement with lattice data for small values of $\hat{\mu}_{B}$. Even for the smaller intensity of magnetic field considered in the lattice simulations of Ref. Borsanyi et al. [2024], our results for the susceptibilities $\chi_{BB}$, $\chi_{QQ}$ and $\chi_{SS}$ are reasonable when compared to their high end in temperatures.
@@
To consider moderate magnetic fields, in a region that is more relevant phenomenologically, one should probably combine perturbative QCD with a model description for lower temperatures and densities. Although this is certainly useful, and a path we plan to follow, one should have in mind that much lower magnetic fields barely affect the equation of state and the phase diagram. Moreover, that would depart from the clean first-principle calculation we have presented in this paper.